A scientist once pointed out that a standard lecture period (50 min) is close to one microcentury.How long is a microcentury in minutes, and what is the percent difference from the scientists approximations?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Shahid
 
1 century = 100 years = 100 * 365.25 days = 36525 days = 36525 *24 *60 min = 52596000 minutes
Hence
1 micro century = 52596000* 10^-6 = 52.596 minutes
Hence
Percentage difference = (52.596 -50)* 100/ 52.596
= 4.9 %
 
Regards
Arun (askIITians forum expert)